OpenStudy (rositaluv):

I have a few square root simplification questions.... Can the square roots of 20, 29, 37,0, 100, 25,and 45? Any help will be greatly appreciated:D

OpenStudy (anonymous):

So for instance, you're looking for like the sqrt(20) = 2sqrt(5) ?

OpenStudy (rositaluv):

I had to use the distance formula and i was told to write it as a square root. I did that. Now I just need some help checking if they can be simplified.

OpenStudy (anonymous):

Look for factors which are perfect squares...for instance 4,8,16,25,etc.

OpenStudy (anonymous):

4,9* my bad

OpenStudy (anonymous):

25 and 100 are perfect squares already

OpenStudy (anonymous):

Oh, no problem. Simply look at the numbers you have here and if you can find two numbers that multiply together to give you the number (one being a perfect square) then it can be simplified. 20: \[2\sqrt{5}\] 29: cannot be simplified 37: cannot be simplified 100: 10*10 = 100 so the answer is 10 25: 5*5 = 25 so the answer is 5 45: \[3\sqrt{5}\]

OpenStudy (rositaluv):

ohh ok i get it.....and i just do that for every one?

OpenStudy (anonymous):

Yup! So like, if you saw 90, you would say "okay 9*10 is 90 and 9 is a perfect square." So then you rewrite your square root as sqrt(9*10) which is 3sqrt(10)

OpenStudy (rositaluv):

so what would i write for 45 again?

OpenStudy (anonymous):

45 = 9*5 so 3sqrt(5)